OpenStudy (anonymous):

An equation is shown below: 2(3x – 5) = 1 Which of the following correctly shows the first two steps to solve this equation? Step 1: 6x – 10 = 1; Step 2: 6x = 11 Step 1: 6x – 5 = 1; Step 2: 6x = 6 Step 1: 5x – 3 = 1; Step 2: 5x = 4 Step 1: 5x – 7 = 1; Step 2: 5x = 8

OpenStudy (anonymous):

step one is to distribute everything out of the parenthesis

OpenStudy (anonymous):

step two is to combine like terms

OpenStudy (anonymous):

so 2(3x - 5) = 1 2*3x=6x 2*-5=-10 6x-10=1 can you do the rest?

OpenStudy (anonymous):

hello? are you there? @agl2029

OpenStudy (anonymous):

yes but i don't know how to do it!

OpenStudy (anonymous):

ok so the next step would be to combine like terms so are since there is only one number with an x attached to it we leave it be. the next combination of terms we can combine are -10 and 1 so now we have: 6x - 10 = 1 so now we have to take the 10 and move it to the opposite side of the equal sign so 6x - 10 = 1 + 10 +10 the tens cancel out on the left side of the equation 6x = 11

OpenStudy (anonymous):

@agl2029 can you come back please?

OpenStudy (anonymous):

so it is the first one of the answer column right?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

ok thanks for the help! it was helfull!

OpenStudy (anonymous):

you are very welcome! :D need any more help? btw can you medal? thnx! :D

OpenStudy (anonymous):

yes i do need more help in pre-algebra! i will close this question and open another on!

OpenStudy (anonymous):

:D